OpenStudy (anonymous):

sec/csc^2 = tan * sin

OpenStudy (bahrom7893):

Sec=1/Cos Csc=1/Sin Tan=Sin/Cos

OpenStudy (anonymous):

(1/cos)(sin^2)=tan sin